OpenStudy (anonymous):

Prove That : ___________ SinA/2=sqrt{(c-b)/2c}

OpenStudy (experimentx):

?? what are b and c??

OpenStudy (anonymous):

variables :D

OpenStudy (anonymous):

look i will tell u where i stopped at :D.. I began with the L.H.S :SinA/2 and then i worked on the law sqrt{(1-cosA)/2}

OpenStudy (asnaseer):

use the relation:\[\sin(A/2)=\frac{1-\cos(A)}{2}\]

OpenStudy (anonymous):

and at the substitution I faced the problem

OpenStudy (asnaseer):

sorry that should be \[\sin^2(A/2)=...\]

OpenStudy (anonymous):

no thats right @asnaseer

OpenStudy (asnaseer):



OpenStudy (experimentx):

is it a right angled triangle??

OpenStudy (asnaseer):

now use this for the right-angled triangle I drew above:\[\sin^2(A/2)=\frac{1-\cos(A)}{2}\]

OpenStudy (anonymous):

Its right angled triangle at C ,

OpenStudy (anonymous):